OpenStudy (anonymous):

find the slope of a function f=3x^2+x-1 when x=2

OpenStudy (mimi_x3):

y' = 6x + 1 y' = 6(2)^2 + 1 m = 25

OpenStudy (saifoo.khan):

c'mon type fast!

OpenStudy (mimi_x3):

me ? it takes time to think man , but im think im wrong xD

OpenStudy (anonymous):

is wrong

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

Yea, u ..

OpenStudy (mimi_x3):

damn , i forgot how do it xD

OpenStudy (mimi_x3):

you do it then saifoo xD

OpenStudy (saifoo.khan):

lol, i also dont remember!

OpenStudy (saifoo.khan):

i did that like 4mnths ago!

OpenStudy (mimi_x3):

lmao , yeah right

OpenStudy (saifoo.khan):

XxD

OpenStudy (mimi_x3):

how is my one wrong ?

OpenStudy (saifoo.khan):

I D K.

OpenStudy (mimi_x3):

There's no other way to find it i think

OpenStudy (anonymous):

6x+1 6*2+1 13

OpenStudy (saifoo.khan):

HAHA, DUMB.

OpenStudy (saifoo.khan):

@MEME.

OpenStudy (saifoo.khan):

Valeriu is Right.

OpenStudy (mimi_x3):

WHAT !? thats not possible , he's like finding y

OpenStudy (saifoo.khan):

lol, idk.

OpenStudy (anonymous):

derivation first

OpenStudy (mimi_x3):

ohh yeah im wrong my bad

OpenStudy (saifoo.khan):

LOL.

OpenStudy (mimi_x3):

i did it half right though

OpenStudy (saifoo.khan):

that wont count, your answer is incorrect.

OpenStudy (mimi_x3):

i dont care , i jst made one silly mistake but i go the derivative right xD

OpenStudy (saifoo.khan):

neither do I. But your answer was wrong.

OpenStudy (mimi_x3):

well , its better than you not knowing how to do it hehe